4/9-2/3+6
I tried 4/9-6/9+6
since 3 can go into 9 I multiplied the numerators by 3..please help I do not know how to get 5 7/9 as the answer.

Answer by Alan3354(69443)

4/9-2/3+6
I tried 4/9-6/9+6
since 3 can go into 9 I multiplied the numerators by 3..please help I do not know how to get 5 7/9 as the answer.
-------------------
4/9-2/3+6
I tried 4/9-6/9+6
= 4/9 - 6/9 + 54/9
= (4 - 6 + 54)/9
= 52/9
= 5 + 7/9